OpenStudy (anonymous):

who wants a medal??????? find the equation of the line having the pair of points (1/6,-1/3) and (5/6, 1) what is the equation of the line?

OpenStudy (tad1):

[-(1/3) - 1]/[( 1/6 )- (5/6} = (-4/3)/(-2/3) = 2= slope y +(1/3) = 2( x + (4/3))